A rectangular gate has a width of 24 inches and a diagonal of 50 inches. How tall is the gate to the nearest inch? I tried to multiply 24 by 2 and 50 by 2, I added that up and divided by 4? I don't think that is correct, please help.
Answer by vihits13(20)

ok the method u used to solve this was completely wrong but thats ok. :-)
you have to use the pythagorean theorem. the pythagorean theorem can only be used on right triangles and it states: a^2 + b^2= c^2.
a and b are the legs of the right triangle and c is the hypotenuse.
in this question, u r given one of the legs and the hypotenuse. the width would be the leg and the diagnol would be the hypotenuse.
lets set up the equation:
24^2+b^2=50^2
576+b^2=2500
b^2=1924 now take the square root of 1924.
b=43.86 the length is obviously b.
the length is approximately equal to 43.86 inches. it would be 44 inches if you want it to the nearest inch.